Write the python function time to midnight() that takes the name of a file containing multiple lines of text giving time in HH:MM:SS HRS or HH:MM:SS AM or HH:MM:SS PM format one per line. The HRS notation indicates CSE 101 – Summer 2019 Lab Assignment #9 2 that the time has been given in 24-hour time format. Each line tells the current time of the day. The function reads the file line by line and calculates the number of seconds remaining until midnight i.e., 00:00:00 HRS. For each line in the file the function calculates the number of seconds remaining until midnight and appends each into a list. The function then returns that list.
Answer #1

def time_to_midnight(): #time to midnight function
   file_name = input("Enter file name: ") #taking file name
   fp1 = open(file_name, 'r') #opening file in read mode
   l=[] #list to store remaining seconds
   for line in fp1: #iterating through each line in file
       time, form = line.split(' ') #splitting every line to time and format based on space
       a =[] #list to store time
       for i in time.split(':'): #splitts time into hours, min, sec
           a.append(int(i))
       form= form.strip() #removing spaces from format
       if (form == "HRS" or form == "AM"): #if the format is HRS or AM then subtract 24-hours and * 3600
           a[0] = (24-a[0])*3600
       elif (form == 'PM'): #if the format is PM then subtract from 12
           a[0] = (12-a[0])*3600
       a[1] = (a[1])*60 #minutes to seconds
       total = (a[0]-a[1])-a[2] #total seconds remaining
       l.append(total) #adding time to list
   print(l) #printing list
  
time_to_midnight()
      


contents in file

ex:

20:15:32 HRS

24 - 20 = 4* 3600 = 14,400 #remaining time to reach midnight

15*60 = 900 #minutes

32 #seconds

sec = 14,400 - 900 = 13,500 - 32 = 13,468.

remaining hours to reach midnight converted to seconds and then the finished minutes and seconds are removed from remaining hours.